OpenStudy (anonymous):

Electrical service costs $8 per month plus $0.10 per kilowatt-hour of electricity used. (Assume that a partial kilowatt-hour is not rounded up.) (a) If the resident of an apartment uses 650 kilowatt hours in 1 month, what is the charg

OpenStudy (anonymous):

.1*650+8=

OpenStudy (anonymous):

73

OpenStudy (texaschic101):

.10(650) + 8 = 65 + 8 = 73

OpenStudy (anonymous):

73